Circular No. 7423

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V1333 AQUILAE R. Jain, C. Bailyn, and P. Coppi, Yale University; M. Garcia, Harvard-

Smithsonian Center for Astrophysics; A. Levine, Massachusetts Institute of Technology; J. Orosz, Utrecht University; P. Lu, Western Connecticut State University; and J. Espinoza and D. Gonzalez, Cerro Tololo Interamerican Observatory (CTIO), report YALO consortium observations: “V1333 Aql = Aquila X-1 appears to be beginning a new optical outburst. This x-ray binary outbursts approximately once per year, and last erupted in May 1999 (IAUC 7161). We have been monitoring this source since Apr. 5 in R using the YALO 1-m telescope at CTIO. Compared to the quiescent average magnitude, which fluctuates by < 0.1 mag, the source had brightened in R by 0.16 ± 0.02 mag on May 11.23 UT. During May 12.23–12.40, we obtained four observations that indicate an increase in brightness of 0.27 to 0.34 ± 0.02 mag compared to the quiescent mean. We will continue monitoring this source at optical, infrared, and x-ray (using RXTE) wavelengths. Observations at all wavelengths are encouraged.”

SUPERNOVA 2000cf IN MCG +11-19-25 S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian Center for As-

trophysics, report that a spectrum of SN 2000cf (cf. IAUC 7421), taken by M. Calkins on May 11.33 UT with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova just past maximum light. The spectrum exhibits a blue continuum with broad Si II, S II, Mg II, Fe II, and Ca II features. Cross-correlation of the host-galaxy spectrum with an emission-line template yields a recession velocity of 10 920 ± 20 km/s for MCG +11-19-25. The supernova expansion velocity, measured with the Si II (rest 635.5 nm) profile, is 9300 km/s. The age of the supernova, based on its spectral features (Riess et al. 1997, A.J. 114, 722), is 5 ± 3 days after maximum light.

SUPERNOVA 2000ce IN UGC 4195 Corrigendum. On IAUC 7422, line 8, for (1591 km/s) yields an ex-

pansion velocity for the supernova of 7600 km/s, read (4888 km/s) yields an expansion velocity for the supernova of 10 900 km/s,

2000 May 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green